A series circuit with a resistor of 100 Ω
capacitor of 25 μF and inductance of
0.15 H is connected across 220-V, 60-Hz
supply. Calculate (i) current (ii) power
and (iii) power factor in the circuit.
1
Expert's answer
2020-09-24T16:01:52-0400

Calculating the impedance of the circuit-

XL = 2pifL = (2)(3.14)(60)(0.15)

= 56.55 ohm

XC = "\frac{1}{2fC\pi}" = "\frac{10^6}{(2)(3.14)(25)(60)}"

= 106.10 ohm

R = 100 ohm

Total Impedance(Z) = (R2 + (XL - XC)2)0.5

Z = (10000 + 2455.2025)0.5

= 111.6 ohm

(i). Current in the circuit(I)-

I = VS/Z

= "\frac{220}{111.6}"

= 1.97 A, 60 Hz

(ii). Power dissipated(P) in series LCR circuit is-

P = "\frac{V^2R}{|Z|^2}"

= "\frac{(48400)(100)}{12455.2025}"

= 388.59 W

(iii). Power factor = R/|Z|

cos"\phi" = "\frac{100}{111.6}"

= 0.896